OpenStudy (anonymous):

A compound has an empirical formula of C4H4O and a molar mass of 204g/mol. What is the molecular formula ?

OpenStudy (doc.brown):

Would it help if I told you the mass of carbon is 12.011g/mol?

OpenStudy (doc.brown):

If there are four moles of carbon, how much does the carbon weigh?

OpenStudy (anonymous):

Try to calculate the weight of something containing 4 C's, 4 H's and 1 O. Perhaps you'll recognize the value you get...

OpenStudy (anonymous):

do what @floppie86 says, then divide 204 by that number to get a factor, x then multiply the number of atoms of each element by x for example, if x=20 (which I'm sure it doesn't), you would multiply C4*20=C80 H4*20=H80 O1*20=O20 so your molecular formula would be C80 H80 O20

OpenStudy (doc.brown):

4 mol C * (12.011g / mol) = 48.044g C 4 mol H * (1.008g / mol) = 4.032g H 1 mol O * (15.999g / mol) = 15.999g O added together C4H4O = 68.075g How many moles of C4H4O do you need to get 204g?